# Continuous Invariance Learning

### v1

@@ 4 Experiments
To evaluate our proposed CIL method, we conduct extensive experiments on two synthetic datasets and three real-world datasets. we compare CIL with 1) Standard Empirical Risk Minimization (ERM), 2) IRMv1 proposed in (Arjovsky et al., 2019), 3) REx in (1) proposed in (Krueger et al., 2021), and 4) GroupDRO proposed in (Sagawa et al., 2019) that minimize the loss of worst group(domains) with increased regularization in training, 5) Diversify proposed in (Lu et al., 2022) and 6) IIBNet proposed in (Li et al., 2022), adding invariant information bottleneck (IIB) penalty in training. For IRMv1, REx, GroupDRO, and IIBNet, we try them on the original continuous domains as well as manually split the dataset with continuous domains into discrete ones (more details in separate subsections below). All the experiments are repeated at least three times and we report the accuracy with standard deviation on each dataset.
@@ 4.1 Synthetic Datasets
4.1.1 Logit Dataset
(a) Spurious correlation of logit dataset
(b) An illustration Yearbook Yao et al. (2022). Images taken from Yao et al. (2022).
Setting. We generate the first synthetic dataset with the invariant feature $\boldsymbol{x}_{v}\in\mathbb{R}^{2}$, spurious feature $\boldsymbol{x}_{s}\in\mathbb{R}^{20}$ and target $\boldsymbol{y}\in\{0,1\}$. The continuous domain $\boldsymbol{t}\in[0,100]$. The conditional distribution of $\boldsymbol{x}_{s}$ given $\boldsymbol{y}$ varies along $\boldsymbol{t}$ as follows:
$\boldsymbol{y}=$$\begin{cases}0,\mbox{ w.p. }0.5,\\ 1,\mbox{ w.p. }0.5,\end{cases}\boldsymbol{x}_{v}\sim$$\begin{cases}\mathcal{N}(y,\sigma^{2}),\mbox{ w.p. }p_{v},\\ \mathcal{N}(-y,\sigma^{2}),\mbox{ w.p. }1-p_{v},\end{cases}\boldsymbol{x}_{s}\sim$$\begin{cases}\mathcal{N}(y,\sigma^{2}),\mbox{ w.p. }p_{s}(\boldsymbol{t}),\\ \mathcal{N}(-y,\sigma^{2}),\mbox{ w.p. }1-p_{s}(\boldsymbol{t}),\end{cases}$
where $p_{v}$ and $p_{s}(\boldsymbol{t})$ are the probabilities of the feature’s agreement with label $\boldsymbol{y}$.
Notably, $p_{s}(\boldsymbol{t})$ varies among domains while $p_{v}$ stays invariant. The observed feature $\boldsymbol{x}$ is a concatenation of $\boldsymbol{x}_{v}$ and $\boldsymbol{x}_{s}$, i.e., $\boldsymbol{x}:=[\boldsymbol{x}_{v},\boldsymbol{x}_{s}]$. We generate 2000 samples as the training dataset $\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})\}_{i=1}^{2000}$. Our goal is to learn a model $f(\boldsymbol{x})$ to predict $\boldsymbol{y}$ that solely relies on $\boldsymbol{x}_{v}$. We set $p_{v}$
to be 0.9 in all domains. $p_{s}(\boldsymbol{t})$ varies across the domains as in Figure 2(a) .
We can see that $\boldsymbol{x}_{s}$ exhibits a high but unstable correlation with $\boldsymbol{y}$. Since existing methods need discrete domains, we equally divide the continuous domain into different numbers of discrete domains, i.e, $\{2,4,8,16\}$.
Env. Type Split Num Method Linear $p_{s}$ Sine $p_{s}$
None ERM 25.33 (2.01) 36.18 (3.06)
Discrete 4 IRMv1 55.39 (6.87) 69.91 (2.40)
8 REx 46.46 (7.50) 68.15 (5.51)
2 GroupDRO 46.68(1.80) 60.95(2.23)
16 IIBNet 51.50(18.60) 49.68(21.87)
Continuous IRMv1 25.90(2.85) 41.28(3.66)
REx 42.73(11.47) 63.33(5.05)
Diversify 53.39(2.96) 53.30(1.93)
CIL 60.95 (6.59) 76.25 (4.99)
Table 2: Comparison on the synthetic Logit data. The metrics including accuracy and standard deviations in parenthesis are calculated on three independent runs. The environment type “Discrete” indicates that we manually create by equally splitting the raw continuous domains. The environment type “Continuous” indicates we use the original continuous domain index. “Split Num” stands for the number of domains we manually create and we report the best performance among spilt $\{2,4,6,8,16\}$. We try settings where the spurious correlation $p_{s}(\boldsymbol{t})$ changes by linear and Sine functions as shown in Fig 2(a), respectively. Detailed results in Appendix L
Results. Table 2 shows the test accuracy of each method (we only present the best split for discrete methods here due to space limit. The details are provided in Appendix K). ERM performs the worst and the large gap implies that ERM models heavily depend on spurious features. IRMv1 improves by 25% on average compared to ERM. CIL outperforms ERM by over 30% on average. CIL also improves by 5-7% over all existing methods based on discrete domains on the best manual discrete domain partition, which indicates that CIL can learn invariant features more effectively. In the appendix, we show the trend of how the performance of IRMv1 and REx changes with the split number.
@@ 4.1.2 Continuous CMNIST
CIL 57.35 (6.89) 57.20 (0689)) 6980 (3.95) 59.50 (8.67)
Table 3: Accuracy on Continuous CMNIST for Linear and Sine $p_{s}(\boldsymbol{t})$. The standard deviation in brackets is calculated with 5 independent runs. The Env. type “Discrete” means that we manually create by equally splitting the raw continuous domains. The environment type “Continuous” indicates using the original continuous domain index. “Split Num” stands for the number of domains we manually create and we report the best performance among spilt $\{2,4,6,8,16\}$. Detailed results in Appendix L
Setting. We construct a continuous variant of CMNIST following (Arjovsky et al., 2019). The digit is the invariant feature $\boldsymbol{x}_{v}$ and the color is the spurious feature $\boldsymbol{x}_{s}$. Our goal is to predict the label of the digit, $\boldsymbol{y}$. We generate 1000 continuous domains. The correlation between $\boldsymbol{x}_{v}$ and the label $\boldsymbol{y}$ is $p_{v}=75\%$, while the spurious correlation $p_{s}(\boldsymbol{t})$ changes among domains $\boldsymbol{t}$, whose details are included in the Appendix L. Similar to the previous dataset, we try two settings with $p_{s}(\boldsymbol{t})$ being a linear and Sine function.
@@ 4.2.1 HousePrice
CIL 82.51(1.96) 79.29 (0.77) 80.30 (2.06) 75.01 (1.18) 81.25 (1.65) 71.29 (0.04)
Table 4: Accuracy of each method on three real-world datasets with standard deviation in brackets. Each method takes 5 runs independently. The details of the settings for HousePrice, Insurance Fraud, and Alipay Auto-scaling can be found in Section 4.2.1, 4.2.2, and 4.2.3, respectively.
@@ 4.2.3 Alipay Auto-scaling
Auto-scaling (Qian et al., 2022) is an effective tool in elastic cloud services that dynamically scale computing resources (CPU, memory) to closely match the ever-changing computing demand. Auto-scaling first tries to predict the CPU utilization based on the current internet traffic. When the predicted CPU utilization exceeds a threshold, Auto-scaling would add CPU computing resources to ensure a good quality of service (QoS) effectively and economically. In this task, we aim to predict the relationship between CPU utilization and the current running parameters of the server in the Alipay cloud. Each record includes 10 related features, such as the number of containers, and network flow, etc. We construct a binary classification task to predict whether CPU utilization is above the $13\%$ threshold or not, which is used in the cloud resource scheduling to stabilize the cloud system. We take the minute of the day as the continuous domain index. The dataset contains 1440 domains with 30 samples in each domain. We then split the continuous domains by consecutive 60 minutes as the discrete domains for IRMv1, REx, GroupDRO, and IIBNet. The data taken between 10:00 and 15:00 are as the testing set because the workload variance in this time period is the largest and it exhibits obviously unstable behavior. All The remaining data serves as the training set. Table 4 reports the performance of all methods. ERM performs the best in training but worst in testing, implying that ERM suffers from the distributional shift. On the other hand, IRMv1 performs the best across all existing methods, exceeding ERM by 9%.
@@ 4.2.4 WildTime-YearBook
We adopt the Yearbook dataset in Wildtime benchmark (Yao et al., 2022)***https://github.com/huaxiuyao/Wild-Time, which is a gender classification task on images taken from American high school students as shown in Figure 2(b). The Yearbook consists of 37K images with the time index as domains. The training set consists of data collected from 1930-1970 and the testing set covers 1970-2013. We adopt the same dataset processing, model architecture, and other settings with the WildTime benchmark. Our method achieves the best OOD performance of 71.22%, improving about 1.5% over the previous SOTA methods (marked with underline) as shown in Table 5.
@@
Table 5: The accuracy on the worst test OOD domain of each method on Yearbook dataset on Wild-time. The performance of baseline methods is copied from Yao et al. (2022).
CIL outperforms ERM and IRMv1 by 13% and 5%, respectively, indicating its capability to recover a more invariant model achieving better OOD generalization.
5 Conclusion
We proposed Continuous Invariance Learning (CIL) that extends invariance learning from discrete categorical indexed domains to natural continuous domain in this paper and theoretically demonstrated that CIL is able to learn invariant features on continuous domains under suitable conditions. Finally, we conducted experiments on both synthetic and real-word datasets to verify the superiority of CIL.
@@ Appendix B Related Work
Causality, invariance, and distribution shift. The invariance property is well known in causal literature back to early 1940s Haavelmo [1944], showing the conditional of the target given its direct causes are invariant under intervention on any node in the causal graph except for the target itself. In 2016, Invariant Causal Prediction (ICP) is first proposed in Peters et al. [2016] to utilize invariance property to identify the direct causes of the target. Models relying on the target’s direct causes are robust to interventions. From the perspective of causality, the distributional shift in the testing distribution is due to the interventions on the causal graph Arjovsky et al. [2019]. So a model with invariance property can hopefully generalize even in the existence of distributional shifts. Anchor regression Rothenhäusler et al. [2018] build the connection between distributional robustness with causality, and demonstrates that a suitable penalty in anchor regression based on the causal structure is equivalent to ensure a certain degree of robustness to distributional shift. Notably, Peters et al. [2016], Rothenhäusler et al. [2018] both assume the inputs $\boldsymbol{x}$ are handcrafted meaningful features, which limits their applications in machine learning and deep learning where the input can be raw images.
@@
Arjovsky et.al. Arjovsky et al. [2019] proposed the first invariance learning method, invariant risk minimization (IRM), which extends ICP Peters et al. [2016] to deep learning by incorporating feature learning. Invariance learning has gained its popularity in recent years and inspired a line of excellent works, where a variety of variants have been proposed. To name a few, Krueger et al. [2021], Xie et al. [2020] penalize the variance of losses in domains; Ahuja et al. [2020] incorporates game theory into invariance learning; Chang et al. [2020] estimates the violation of invariance by training multiple independent networks; Jin et al. [2020] proposes an invariance penalty based on regret. Notably, these works all require discretely indexed domains. Another line of works try to learn invariant features when explicit domain indices are not provided Creager et al. [2021]. However, Lin et al. [2022b] theoretically shows that it is generally impossible to learn invariance without environmental information. Some recent works Lin et al. [2022a], Zhou et al. [2022] show that invariance learning methods are sensitive to overfitting caused by the overparameterization of deep neural networks. These works are orthogonal to our study.
@@
Distribution Shift with Continuous Domains. There are a few methods considering domain shifts with continuous domain indexes. Bobu et al. [2018], Hoffman et al. [2014], Wulfmeier et al. [2018], Bitarafan et al. [2016] consider the distribution shifts incrementally over time, while trying to perform domain adaptation sequentially. Continuously indexed domain adaption (CIDA) Wang et al. [2020] is the first to adapt across (multiple) continuously indexed domains simultaneously. Note that these works all assume that the input $\boldsymbol{x}$ of the samples in the testing domains are available, and are therefore not applicable for the OOD generalization tasks considered in this paper. Zhang et al. [2017b] tries to discover causal graph based on the continuous heterogeneity, but assumes that the features are given (cannot be learned from raw input), making it not applicable in our setting either.
@@
Invariance Property in Causality. Figure 3 shows an example (similar to Figure 1 of ICP Peters et al. [2016]) of an causal system with nodes $(y,x_{1},x_{2},x_{3},x_{4})$. Suppose our task is to build a model based on a subset of $\{x_{1},x_{2},x_{3},x_{4}\}$ to predict $y$, under the distributional shifts in different domains caused by interventions. In our example, there are interventions on node $x_{2}$ in domain 2, and interventions on nodes $x_{3}$ and $x_{4}$ in domain 3. Notably, we do not allow interventions on target $y$ itself (indicating that the noise ratio of $y$ should be constant among domains). The invariance property shows that the conditional probability of $y$ given its parents remains the same on interventions on any nodes except for the $y$ itself. In this example, $P(y|x_{2},x_{3})$ remains the same under all the three domains. However, one can check that $P(y|x_{1})$ and $P(y|x_{4})$ changes in domain 2 and 3, respectively. Therefore it is safe to build the model on $(x_{2},x_{3})$ to predict $y$. In contrast, $x_{1}$ and $x_{4}$ are unreliable because the conditional distribution is unstable under distributional shifts. In this example, $\boldsymbol{x}_{v}$ is $\{x_{2},x_{3}\}$, and $\boldsymbol{x}_{s}$ is $\{x_{1},x_{4}\}$.
Figure 3: An illustration of the invariance property in causality (similar to Figure 1 in ICP Peters et al. [2016]). This figure shows a causal system with five nodes, $y$, $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$. Our task is to predict $y$ based on the $x^{\prime}$s. There are different interventions in different domains, leading to distributional shifts. Intervention on a node can be simply interpreted as changing the node value. The changes can propagate to the descendants of the intervened node. The invariance property shows that $P(y|x_{2},x_{3})$ remains the same in all three domains. In contrast, $P(y|x_{1})$ and $P(y|x_{4})$ changes in domain 2 and 3 due to interventions, respectively. So it is safe to build model on $x_{2}$ and $x_{3}$ to predict $y$, which is expected to be stable under novel testing distribution.

### v2

@@ 4 Experiments
To evaluate our proposed CIL method, we conduct extensive experiments on two synthetic datasets and four real-world datasets, the synthetic logit dataset is presented in Appendix K. we compare CIL with 1) Standard Empirical Risk Minimization (ERM), 2) IRMv1 proposed in Arjovsky et al. (2019), 3) REx in equation 1 proposed in Krueger et al. (2021), and 4) GroupDRO proposed in Sagawa et al. (2019) that minimize the loss of worst group(domains) with increased regularization in training, 5) Diversify proposed in Lu et al. (2022) and 6) IIBNet proposed in Li et al. (2022), adding invariant information bottleneck (IIB) penalty in training. Note that while CIDA (Wang et al., 2020) and its variants (Xu et al., 2022; 2023; Liu et al., 2023) also handle continuously indexed domains, they are domain adaptation methods and therefore not included as baselines (see Appendix B for details). For IRMv1, REx, GroupDRO, and IIBNet, we try them on the original continuous domains as well as manually split the dataset with continuous domains into discrete ones (more details in separate subsections below). All the experiments are repeated at least three times and we report the accuracy with standard deviation on each dataset.
@@ 4.1.1 Continuous CMNIST
CIL (Ours) 57.35 (6.89) 57.20 (6.89) 69.80 (3.95) 59.50 (8.67)
Table 2: Accuracy on Continuous CMNIST for Linear and Sine $p_{s}(\bm{t})$. The standard deviation in brackets is calculated with 5 independent runs. The Env. type “Discrete” means that we manually create by equally splitting the raw continuous domains. The environment type “Continuous” indicates using the original continuous domain index. “Split Num” stands for the number of domains we manually create and we report the best performance among spilt $\{2,4,8,16\}$. Detailed results in Appendix L
Setting. We construct a continuous variant of CMNIST following Arjovsky et al. (2019). The digit is the invariant feature $\bm{x}_{v}$ and the color is the spurious feature $\bm{x}_{s}$. Our goal is to predict the label of the digit, $\bm{y}$. We generate 1000 continuous domains. The correlation between $\bm{x}_{v}$ and the label $\bm{y}$ is $p_{v}=75\%$, while the spurious correlation $p_{s}(\bm{t})$ changes among domains $\bm{t}$, whose details are included in the Appendix L. Similar to the previous dataset, we try two settings with $p_{s}(\bm{t})$ being a linear and Sine function.
@@ 4.2.1 HousePrice
IIBNet 62.29 (4.40) 53.93 (3.70) 76.34 (5.20) 72.01 (6.99) 80.49 (8.30) 58.89 (5.53)
InvRAT 82.29 (0.76) 77.18 (0.41) 80.63 (1.04) 72.07 (0.74) 88.74 (1.54) 60.58 (3.22)
EIIL 82.62 (0.42) 76.85 (0.44) 80.60 (1.36) 72.44 (0.58) 91.34 (1.50) 53.14 (0.74)
HRM 84.67 (0.62) 77.40 (0.27) 81.93 (1.11) 73.52 (0.46) 89.84 (1.20) 55.44 (0.35)
ZIN 84.80 (0.60) 77.54 (0.30) 81.93 (0.73) 73.33 (0.43) 90.56 (0.91) 58.99 (0.87)
CIL (L1) 83.41 (0.75) 77.98 (1.02) 82.39 (1.40) 76.54 (1.03) 81.44 (1.77) 68.51 (1.33)
CIL (L2) 82.51 (1.96) 79.29 (0.77) 80.30 (2.06) 75.01 (1.18) 81.25 (1.65) 71.29 (0.04)
Table 3: Accuracy of each method on three real-world datasets with standard deviation in brackets. Each method takes 5 runs independently. The details of the settings for HousePrice, Insurance Fraud, and Alipay Auto-scaling can be found in Section 4.2.1, 4.2.2, and 4.2.3, respectively. CIL is our method. L1 or L2 means we use the L1 or L2 loss. We adopt L2 loss by default in other tables.
@@ 4.2.3 Alipay Auto-scaling
Auto-scaling (Qian et al., 2022) is an effective tool in elastic cloud services that dynamically scale computing resources (CPU, memory) to closely match the ever-changing computing demand. Auto-scaling first tries to predict the CPU utilization based on the current internet traffic. When the predicted CPU utilization exceeds a threshold, Auto-scaling would add CPU computing resources to ensure a good quality of service (QoS) effectively and economically. In this task, we aim to predict the relationship between CPU utilization and the current running parameters of the server in the Alipay cloud. Each record includes 10 related features, such as the number of containers, and network flow, etc. We construct a binary classification task to predict whether CPU utilization is above the $13\%$ threshold or not, which is used in the cloud resource scheduling to stabilize the cloud system. We take the minute of the day as the continuous domain index. The dataset contains 1440 domains with 30 samples in each domain. We then split the continuous domains by consecutive 60 minutes as the discrete domains for IRMv1, REx, GroupDRO, and IIBNet. The data taken between 10:00 and 15:00 are as the testing set because the workload variance in this time period is the largest and it exhibits obviously unstable behavior. All the remaining data serves as the training set. Table 3 reports the performance of all methods. ERM performs the best in training but worst in testing, implying that ERM suffers from the distributional shift. On the other hand, IRMv1 performs the best across all existing methods, exceeding ERM by 9%. CIL outperforms ERM and IRMv1 by 13% and 5%, respectively, indicating its capability to recover a more invariant model achieving better OOD generalization.
@@ 4.2.4 WildTime-YearBook
We adopt the Yearbook dataset in Wildtime benchmark (Yao et al., 2022)***https://github.com/huaxiuyao/Wild-Time, which is a gender classification task on images taken from American high school students as shown in Figure 4. The Yearbook consists of 37K images with the time index as domains. The training set consists of data collected from 1930-1970 and the testing set covers 1970-2013. We adopt the same dataset processing, model architecture, and other settings with the WildTime. To keep consistent with Yao et al. (2022), we adopt the same baseline methods with Yao et al. (2022) and directly copy the performance of baseline methods from Yao et al. (2022).For the details of these baselines, we refer the reader to the Appendix B of Yao et al. (2022). Table 4 shows that we achieve the best OOD performance of 71.22%, improving about 1.5% over the previous SOTA methods (marked with underline).
Figure 3: An illustration Yearbook (Yao et al., 2022). Images taken from Yao et al. (2022).
@@
Figure 4: The accuracy on the worst test OOD domain of each method on Yearbook dataset on Wild-time. The performance of baseline methods is copied from Yao et al. (2022).
5 Conclusion and Discussion
We proposed Continuous Invariance Learning (CIL) that extends invariance learning from discrete categorical indexed domains to natural continuous domain in this paper and theoretically demonstrated that CIL is able to learn invariant features on continuous domains under suitable conditions. However, learning invariance would be more challenging with larger DNNs due to IRM’s inherent sensitivity to over-fitting (Lin et al., 2022a; Zhou et al., 2022). Recent works has shown the effectiveness of so called spurious feature diversification (Lin et al., 2023a), which has shown very promising performance even on modern large language models (Lin et al., 2023b). It would be an interesting future direction to explore feature diversification on continuous domains.
@@ Appendix B Related Work
Causality, invariance, and distribution shift. The invariance property is well known in causal literature back to early 1940s (Haavelmo, 1944), showing the conditional of the target given its direct causes are invariant under intervention on any node in the causal graph except for the target itself. In 2016, Invariant Causal Prediction (ICP) is first proposed in Peters et al. (2016) to utilize invariance property to identify the direct causes of the target. Models relying on the target’s direct causes are robust to interventions. From the perspective of causality, the distributional shift in the testing distribution is due to the interventions on the causal graph (Arjovsky et al., 2019). So a model with invariance property can hopefully generalize even in the existence of distributional shifts. Anchor regression (Rothenhäusler et al., 2018) build the connection between distributional robustness with causality, and demonstrates that a suitable penalty in anchor regression based on the causal structure is equivalent to ensure a certain degree of robustness to distributional shift. Notably, Peters et al. (2016); Rothenhäusler et al. (2018) both assume the inputs $\bm{x}$ are handcrafted meaningful features, which limits their applications in machine learning and deep learning where the input can be raw images.
@@
Arjovsky et al. (2019) proposed the first invariance learning method, invariant risk minimization (IRM), which extends ICP (Peters et al., 2016) to deep learning by incorporating feature learning. Invariance learning has gained its popularity in recent years and inspired a line of excellent works, where a variety of variants have been proposed. To name a few, Krueger et al. (2021); Xie et al. (2020) penalize the variance of losses in domains; Ahuja et al. (2020) incorporates game theory into invariance learning; Chang et al. (2020) estimates the violation of invariance by training multiple independent networks; Jin et al. (2020) proposes an invariance penalty based on regret. Notably, these works all require discretely indexed domains. Another line of works try to learn invariant features when explicit domain indices are not provided (Creager et al., 2021). However, Lin et al. (2022b) theoretically shows that it is generally impossible to learn invariance without environmental information. Some recent works (Lin et al., 2022a; Zhou et al., 2022) show that invariance learning methods are sensitive to overfitting caused by the overparameterization of deep neural networks. These works are orthogonal to our study.
@@
Distribution Shift with Continuous Domains. There are a few methods considering domain shifts with continuous domain indexes. Bobu et al. (2018); Hoffman et al. (2014); Wulfmeier et al. (2018); Bitarafan et al. (2016) consider the distribution shifts incrementally over time, while trying to perform domain adaptation sequentially. Continuously indexed domain adaption (CIDA) (Wang et al., 2020) is the first to adapt across (multiple) continuously indexed domains simultaneously. Note that these works all assume that the input $\bm{x}$ of the samples in the testing domains are available, and are therefore not applicable for the OOD generalization tasks considered in this paper. Zhang et al. (2017b) tries to discover causal graph based on the continuous heterogeneity, but assumes that the features are given (cannot be learned from raw input), making it not applicable in our setting either.
@@
Figure 5 shows an example (similar to Figure 1 of ICP (Peters et al., 2016)) of an causal system with nodes $(y,x_{1},x_{2},x_{3},x_{4})$. Suppose our task is to build a model based on a subset of $\{x_{1},x_{2},x_{3},x_{4}\}$ to predict $y$, under the distributional shifts in different domains caused by interventions. In our example, there are interventions on node $x_{2}$ in domain 2, and interventions on nodes $x_{3}$ and $x_{4}$ in domain 3. Notably, we do not allow interventions on target $y$ itself (indicating that the noise ratio of $y$ should be constant among domains). The invariance property shows that the conditional probability of $y$ given its parents remains the same on interventions on any nodes except for the $y$ itself. In this example, $P(y|x_{2},x_{3})$ remains the same under all the three domains. However, one can check that $P(y|x_{1})$ and $P(y|x_{4})$ changes in domain 2 and 3, respectively. Therefore it is safe to build the model on $(x_{2},x_{3})$ to predict $y$. In contrast, $x_{1}$ and $x_{4}$ are unreliable because the conditional distribution is unstable under distributional shifts. In this example, $\bm{x}_{v}$ is $\{x_{2},x_{3}\}$, and $\bm{x}_{s}$ is $\{x_{1},x_{4}\}$.
Figure 5: An illustration of the invariance property in causality (similar to Figure 1 in ICP (Peters et al., 2016)). This figure shows a causal system with five nodes, $y$, $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$. Our task is to predict $y$ based on the $x^{\prime}$s. There are different interventions in different domains, leading to distributional shifts. Intervention on a node can be simply interpreted as changing the node value. The changes can propagate to the descendants of the intervened node. The invariance property shows that $P(y|x_{2},x_{3})$ remains the same in all three domains. In contrast, $P(y|x_{1})$ and $P(y|x_{4})$ changes in domain 2 and 3 due to interventions, respectively. So it is safe to build model on $x_{2}$ and $x_{3}$ to predict $y$, which is expected to be stable under novel testing distribution.